OpenStudy (anonymous):

SAT subject test problem: what is the value of k if... (I will post the equation below )

OpenStudy (anonymous):

\[\frac{ 1 }{ (x-2)(x+4) }=\frac{ h }{ (x-2) }+\frac{ k }{ (x+4) }\]

OpenStudy (anonymous):

multiply both sides by (x-2)(x+4) 1 = h*(x+4)+k*(x-2) subtract the h*(x+4) to the other side 1-h*(x+4) = k*(x-2) divide both sides by x-2 (1-h*(x+4))/(x-2) = k

OpenStudy (anonymous):

@tomo but that is not one of the answer choices. here are the choices: a) 2 b)-4 c)-1/6 d)1/6 e)-2

OpenStudy (anonymous):

unless you have a initial value for x and h, i don't now how to get a number for k.

OpenStudy (anonymous):

oh, ok. Thank you